OpenStudy (anonymous):

how to complete the square x^2+9x=-14

OpenStudy (solomonzelman):

\[x^2+9x+20.25=-14+20.25\]

OpenStudy (anonymous):

why 20.25?

OpenStudy (ddcamp):

\[20.25=(\frac{ 9 }{ 2 })^2\]

OpenStudy (solomonzelman):

\[x^2+9x+20.25=6.25\]\[(x+4.5)^2=6.25\]\[(x+4.5)=±~2.5\]\[x=-4.5±~2.5\] Why 20.25 because 4.5^2 or the last term when you would expand the parenthesis.

OpenStudy (anonymous):

i dont understand how to get the 20.25.... or the 9/2 ^2

OpenStudy (solomonzelman):

\[x^2+9x=-14\]what do you add to both sides to make it a perfect square?

OpenStudy (solomonzelman):

you should get \[(x+4.5)^2\]becuase the middle term is then going to be 9 and the first term x^2, but however the last term would be 4.5^2 or 20.25, so this is the number - 20.25 that you should add to both sides, because it's the number that makes the left side into a perfect square.